OpenStudy (anonymous):

solve for x 3x + 10x =15

OpenStudy (anonymous):

x=15/13

OpenStudy (mary.rojas):

add up the x's first

OpenStudy (mary.rojas):

then get the x by itself. 13x=15 so divide the 13 to both sides it will cancel out on 13x side so the x will be by itself then 15 will turn to 15/13 so....x=15/13